A raffle is held and the grand prize is a $1000 plasma television. The tickets are sold for $1 each, if 1000 tickets were sold, find the expected value of one ticket.
1
Expert's answer
2020-09-15T12:17:28-0400

The expected value of one ticket is given by,


The probability of winning the prize * net prize value + The probability of losing the prize * cost of one ticket


"= \frac {1}{1000}*(1000-1)+\frac{999}{1000}*(-1)=\frac {999}{1000}-\frac {999}{1000}=0"


Hence the expected value of one ticket = $0